Consider the following gasoline sales time series. If needed, round your answers to two-decimal digits. Week Sales (1,000s of business analytics
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

The exponential smoothing (ES) forecast with smoothing constant a = 0.1 for the nth period is computed using the following fo

ES(0.1) Forecasts 17 17 +0.1. (17 – 17) = 17 17 +0.1. (21 – 17) = 17.4 17.4 +0.1. (16 – 17.4) = 17.26 17.26 +0.1. (24 - 17.26

Data ES(0.2) Forecasts 17 17+0.2. (17 – 17) = 17 17+0.2. (21 – 17) = 17.8 17.8 +0.2. (16 - 17.8) = 17.44 17.44 +0.2. (24 - 17

Data ES(0.1) Forecasts Abs. Error Error2 Abs. % Error 17 - 17 = 0 (17 – 17) = 0 17-19 = 0% 21 - 17 = 21 - 17) = 16 19.05% = 1

Data ES(0.2) Forecasts Abs. Error Error2 Abs. % Error 17 - 17 = 17 - 17) = 0 17-19 = 0% 0 21 - 17 = (21 - 17) = 16 19.05% 16

Answer a) Thus, Week 13 Forecast for a = 0.1 is 19.1109 Thus, Week 13 Forecast for a = 0.2 is 20.0254 Answer b) Based on erro